OpenStudy (anonymous):

Solve. If there are X teams in a sports league and all the teams play each other twice, a total of N(x) gamers are played, where N(x)=x^2-x. A soccer league has 10 teams, each of which plays the others twice. If the league pays $46 per game for the field and officials, how much will it cost to play the entire schedule?

OpenStudy (anonymous):

\[N(10)=10^2-10=100-10-90\]

OpenStudy (anonymous):

scuse me i meant \[N(10)=10^2-10=90\]

OpenStudy (anonymous):

number of games is 90, so cost is \[90\times 46=4140\]

OpenStudy (anonymous):

ok. i see thanks. medal